OpenStudy (anonymous):

ALGEBRA 1 HELP! * I've tried putting this into slope intercept form using point slope formula, but it doesn't seem to make sense.* Omar is on a diet to lose some weight. He is losing weight at a rate of 2 pounds per week. After 6 weeks, his weight is 204 pounds. Write and solve a linear equation to find how many weeks it will take Omar to reach his target wait of 180 pounds.

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

I got an ordered pair of 6,204

OpenStudy (anonymous):

Then i plugged the numbers in

OpenStudy (anonymous):

But when I tried to solve for 180 pounds it didn't add up

OpenStudy (anonymous):

yes \((6,204)\) would be a point on the line what is the slope?

OpenStudy (anonymous):

m=2

OpenStudy (anonymous):

what did you "plug the numbers in " to

OpenStudy (anonymous):

y-204=2(x-6)

OpenStudy (anonymous):

oooh lol, he is not gaining two pounds a week, he is losing two pounds a week!

OpenStudy (anonymous):

your slope is making him fatter and fatter no wonder you can't solve it

OpenStudy (anonymous):

so is it negative 2?

OpenStudy (anonymous):

yes, he is losing, not gaining

OpenStudy (anonymous):

OHH ok Let me try it

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@satellite73 I got y=-2+192

OpenStudy (anonymous):

Then I subsituted the variables

OpenStudy (anonymous):

and got x=6

OpenStudy (anonymous):

so is it 12 weeks?

OpenStudy (anonymous):

you want me to check?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

there is a mistake somewhere

OpenStudy (anonymous):

in the word problem?

OpenStudy (anonymous):

or was it me let me double check

OpenStudy (anonymous):

\[y-204=-2(x-6)\] distribute get \[y-204=-2x+12\] i bet that was the mistake

OpenStudy (anonymous):

add and get \[y=-2x+216\]

OpenStudy (texaschic101):

y - y1 = m(x - x1) (6,204)....x1 = 6 and y1 = 204 slope(m) = 2 y - 204 = -2(x - 6) y = -2x + 12 + 204 y = -2x + 216 180 = -2x + 216 180 - 216 = -2x -36 = -2x 18 = x not sure if I did that right...but I think your supposed to sub in 180 for y

OpenStudy (anonymous):

ohhhh ok thank you both

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I got it added wrong